On January 1,2020, Jhona Co. acquired 25% interest in the ordinary shares of Eldon Inc. for ₱3,000,000 which reflected book value as of the date. The equipment has a remaining life of five years.


2020 2021

Net income ₱1,000,000 ₱1,500,000

Dividend declared 400,000 700,000

Questions:

Case No. 1: On January 3,2020 Jhona Company sold an equipment costing ₱600,000 to Eldon Company for ₱800,000.

1. Investment income 2020

   2. Investment 2021

Case No. 2: On January 3,2020 Eldon Company sold an equipment costing ₱400,000 to Jhona Company in 2020.

3.Investment income in 2020

4. Investment income 2021



1
Expert's answer
2021-06-17T09:47:15-0400

Solution:

Case No. 1:

1.). Investment income 2020:

Net income 2020 = 1,000,000

Share of net income = 1,000,000 "\times" 25"\%" = 250,000

Adjust for the equipment sale:

Profit from sale of equipment = 800,000 – 600,000 = 200,000

Share of profit from equipment sale = 200,000 "\times" 25"\%" = 50,000

Add to the share of net income = 250,000 + 50,000 = 300,000

Investment income in 2020 = 300,000


Investment income in 2021:

Net income 2021 = 1,500,000

Share of net income = 1,500,000 "\times" 25"\%" = 375,000

Investment income in 2021 = 375,000


Case No. 2:

1.). Investment income in 2020:

Net income 2020 = 1,000,000

Share of net income = 1,000,000 "\times" 25"\%" = 250,000

Adjust for the equipment sale:

Sale of equipment = 400,000

Share of profit from equipment sale = "\frac{400,000}{5} \times 25\%" = 20,000

Deduct from the share of net income = 250,000 - 20,000 = 230,000

Investment income in 2020 = 230,000


Investment income in 2021:

Net income 2021 = 1,500,000

Share of net income = 1,500,000 "\times" 25"\%" = 375,000

Investment income in 2021 = 375,000